An ion with a charge of +3.2 × 10–19 C and a mass of 2.7 × 10–26 kg is moving due south at a speed of 4.8 × 103 m s–1 . It enters a uniform magnetic !eld of strength 4.6 × 10–4 T directed downwards towards the ground. Determine the force acting on the ion. 


1
Expert's answer
2021-12-29T12:35:15-0500

"F=qvB\sin \theta,"

"\theta=90°,~\sin\theta=1,"

"F=qvB=3.2\cdot 10^{-19}\cdot 4.8\cdot 10^3\cdot 4.6\cdot 10^{-4}=7.1\cdot 10^{-21}~N=7.1~zN."